OpenStudy (anonymous):

how do make a table for y= -3x² using y=ax² x: y:

OpenStudy (anonymous):

plug in any values of x and you will get y

OpenStudy (anonymous):

case1:a=-3 x:(-\[\infty\],\[\infty\]) y:(0,\[\infty\]) case2:a\[\neq\]-3 x:0 y:0